OpenStudy (anonymous):

Find the value of a if the lines ax+3y=5 and 2x+(a+1)y=6 are parallel.

OpenStudy (anonymous):

parallel means equal slopes : 3y = 5 - ax y = (5/3)- (ax/3) so slope -a/3 2x + (a+1)y = 6 (a+1)y = 6 - 2x y = (6/(a+1)) - 2x/(a+1) so slope -2/(a+1) now you have to equate them : -2/(a+1) = -a/3 6 = a(a+1) a^2 + a -6 = 0 a^2 +3a -2a -6 = 0 a(a+3) -2 (a+3) (a-2)(a+3) = 0 a = 2, -3

OpenStudy (anonymous):

Thanks :) i forgot a step

OpenStudy (anonymous):

How would you mathematically find the perpendicular?